OpenStudy (anonymous):

Ryan and Ming are selling cookie dough for a school fundraiser. Customers can buy packages of white chocolate chip cookie dough and pacakges of gingerbread cookie dough. Ryan sold 13 packages of white chocolate chip cookie dough and 14 packages of gingerbread cookie dough for a total of $244. Ming sold 6 packages of white chocolate chip cookie dough and 4 packages of gingerbread cookie dough for a total of $88. What is the cost each of one package of white chocolate chip cookie dough and one package of gingerbread cookie dough?

OpenStudy (anonymous):

13x+14y=244 and 6x+4y=88 @phi

OpenStudy (anonymous):

the second equation oes to 3x+2y=44

OpenStudy (anonymous):

we have to figure out how to get the number before x the same on both equations right?

OpenStudy (anonymous):

can't we do that easier with y though?

OpenStudy (phi):

yes, pick the easier one

OpenStudy (anonymous):

hmm I'm trying to think of the least common denominator (that's what it's called right?)

OpenStudy (anonymous):

can't we just do 2 times 7 to get the y in the second equation to 14 as well?

OpenStudy (anonymous):

or do you have to do something to the first equation too?

OpenStudy (anonymous):

I hope I'm making sense haha

OpenStudy (phi):

least common multiple yes 14 is what you want

OpenStudy (anonymous):

okay cool

OpenStudy (anonymous):

21x+14y=308

OpenStudy (phi):

technically you could multiply the top equation by 1 if you want. (but that is being a bit silly)

OpenStudy (phi):

ok, now you have 13x+14y=244 21x+14y=308

OpenStudy (anonymous):

x=8 ?

OpenStudy (anonymous):

y=10?

OpenStudy (phi):

yes if we subtract bottom from top (though I would do it the other way round) , we get -8x = -64 x = -64/-8 = 8

OpenStudy (anonymous):

great :) thank you for all your help!!

OpenStudy (phi):

yw